OpenStudy (anonymous):

simplify: (a^2b^8c^5)^2(a^5c^2)^3(a^-2bc^-1)^3

OpenStudy (anonymous):

LOL mountain prob.

OpenStudy (apoorvk):

 Now multiply the powers outside the brackets, into the powers of each term inside the bracket. and use these fundae: \[(a^m)^n = a^{mn}\] \[a^m \times a^n = a ^{m+n}\]